OpenStudy (anonymous):

Find the thex-coordinate of the vertex of the function given below. Y=(x-5)(x+3)

OpenStudy (anonymous):

The easiest way to determine the coordinates of the vertex is to write the parabola in vertex form.

OpenStudy (anonymous):

Meaning, \[y = a(x-h)^2+k\]

OpenStudy (anonymous):

You will have to expand that expression, then complete the square to get the vertex form.

OpenStudy (anonymous):

You will be able to read off the coordinates of the vertex (h, k) in the vertex form.

OpenStudy (anonymous):

What do I fill in?? For the problem?

OpenStudy (anonymous):

I listed the steps above. First, expand the expression. Then, complete the square.

OpenStudy (anonymous):

I don't know what that means I need to find the a,h,x,k

OpenStudy (anonymous):

Can you first expand the expression?

OpenStudy (anonymous):

What does that mean??

OpenStudy (anonymous):

(x+1)(x+2) = x^2 +3x + 2

OpenStudy (anonymous):

Expand the expression you have, like the example above.

OpenStudy (anonymous):

X^2-2x+3??

OpenStudy (anonymous):

Check your constant term again.

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

You multiply (x-5) by (x+3) right? You should get -5*3 = -15.

OpenStudy (anonymous):

\[Y = (x-5)(x+3) = x^2-2x -15\]

OpenStudy (anonymous):

Do you agree?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now, complete the square.

OpenStudy (anonymous):

This means you need to factorize the expression so that you force it to become a square (x+a)^2

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[x^2-2x-15= (x-1)^2+k\]

OpenStudy (anonymous):

Do you follow what I did in that last step?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I was able to factorize it into (x-1)^2 because of the -2x term.

OpenStudy (anonymous):

But I will get a constant k.

OpenStudy (anonymous):

x^2+1+k?

OpenStudy (anonymous):

No, it's (x-1)^2+k

OpenStudy (anonymous):

Since you're only interested in the x-coordinate of the vertex, we are done. You don't need to solve for k in this question.

OpenStudy (anonymous):

Match the equation that we have, with the vertex formula above.

OpenStudy (anonymous):

What is the value of h?

OpenStudy (anonymous):

\[Y = a(x-h)^2+k =(x-1)^2+k\]

OpenStudy (anonymous):

so what do I do??

OpenStudy (anonymous):

Tell me the value of h.

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Correct, that's your x-coordinate.

OpenStudy (anonymous):

So the answer is 1